POLARIZATION SPECTROSCOPY OF LITHIUM DIMERS IN A^{1}Σ_{u}^{+}-X^{1}Σ_{u}^{+} TRANSITION \underline{X. WANG}, J. YANG, J. QI and A. M. LYYRA, Department of Physics, Temple University, Philadelphia, PA.